OpenStudy (anonymous):

Evaluate. ^3 sqrt -64 -4 4 undefined

OpenStudy (solomonzelman):

cube root (like any nth root, for an odd integer n) is never undefined.

OpenStudy (solomonzelman):

only EVENth root can result in 'i' but ODDth root will not.

OpenStudy (solomonzelman):

now, hint: 4^3=64

OpenStudy (anonymous):

so its 4?

OpenStudy (jdoe0001):

well... let's see 4 * 4 * 4 = -64 ?

OpenStudy (jdoe0001):

\(\large \bf \sqrt[3]{-64}\implies \sqrt[3]{-1\cdot 64}\implies \sqrt[3]{-1}\cdot \sqrt[3]{64}\implies \sqrt[3]{(-1)^3}\cdot \sqrt[3]{64} \)

OpenStudy (solomonzelman):

you are close, (-4)^3=-64

OpenStudy (solomonzelman):

(you is you, and jdoe is correct. Not that it takes me to verify that)

OpenStudy (anonymous):

so -4

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

One more please, can you tell me what will \(\large\color{black}{ \displaystyle \sqrt[3]{-8} }\) be?

OpenStudy (anonymous):

2 ^3 sqrt -1?

OpenStudy (solomonzelman):

(-2)^3=(-8) just as 2^3=8

OpenStudy (solomonzelman):

so, \(\large\color{slate}{\displaystyle\sqrt[3]{-8}=-2}\)

OpenStudy (anonymous):

ohh ok thnx